OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. (5x - 1)/2 + (-3x - 1)/4 = 9/2

OpenStudy (m):

2(5x-1) + (-3x - 1) = (4)(9/2) 10x - 1 + 3x - 1 = 36/2 13x - 2 = 18 13x = 20 x = 20/13

OpenStudy (m):

whoops mistake

OpenStudy (m):

2(5x-1) + (-3x - 1) = (4)(9/2) 10x - 2 + 3x - 1 = 36/2 13x - 3 = 18 13x = 21 x = 21/13

OpenStudy (anonymous):

heyyyy

OpenStudy (m):

whoops mistake again!!!

OpenStudy (anonymous):

lol heyyy haha and omg M!!!! LOL

OpenStudy (radar):

(5x-1)2 + (-3x - 1)/4 =9/2 I would multiply all terms by 4 2(5x-1) +(-3x - 1) =18 remove parens 10x-2-3x-1=18 Now consolidate similar terms. 7x-3=18 add 3 to both sides 7x=21 x= 3 dividing both sides by 7

OpenStudy (anonymous):

whats up

OpenStudy (m):

yeah -3x is correct

OpenStudy (anonymous):

nm my dad is setting up a new computer for me and im finishing up my hm :P

OpenStudy (m):

learn from my common mistakes :)

OpenStudy (anonymous):

alrighty!! thanks guys wow those are like not the steps i was using ^^ and yea i do that all the time dont worry!

OpenStudy (anonymous):

i did all my other problems and now im doing the ones that i couldnt do >.< prolly gonna be posting like a few more questions ^^